soybean meal is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb mixture that is 13% protein?
Answer by checkley77(12844)

.18X+.09(360-X)=.13*360
.18X+32.4-.09X=46.8
.09X=46.8-32.4
.09X=14.4
X=14.4/.09
X=160 POUNDS OF SOYBEAN IS NEEDED.
360-160=200 POUNDS OF CORNMEAL IS NEEDED.
PROOF:
.18*160+.09*200=.13*360
28.8+18=46.8
46.8=46.8